merald Marine Stores Company manufactures decorative fittings for luxury yachts that require highly skilled labor, and special metallic materials.


Emerald uses standard costs to prepare its flexible budget. For the first quarter of 2015, direct material and direct labor standards for one of their popular products were as follows:

Direct materials: 1 pound per unit; $12 per pound
Direct labor: 4 hours per unit; $15 per hour

Emerald produced 5,000 units during the quarter. At the end of the quarter, an examination of the materials records showed that the company used 7,000 pounds of materials and actual total material costs were $98,000.

Calculate the direct materials PRICE Variance.

Group of answer choices

$38,000 Unfavorable

$38,000 Favorable

$14,000 Unfavorable

$14,000 Favorable
Answer #1

Materials Price Variance

= (SP-AP) *AQ

= (12 - 98,000/7,000)*7,000

= (12 - 14) * 7,000

= 14,000 UNFAVORABLE